Determine the complex number Z which satisfies the equations |z+3i|=|z+5-2i| and |z-4i|=|z+2i| simultaneously
1
Expert's answer
2020-05-19T19:14:55-0400

Let "z=a+bi." Then


"|z+3i|=|z+5-2i|""|a+(b+3)i|=|(a+5)+(b-2)i|"

"a^2+(b+3)^2=(a+5)^2+(b-2)^2""a^2+b^2+6b+9=a^2+10a+25+b^2-4b+4""10a=10b-20""a=b-2"


"|z-4i|=|z+2i|""|a+(b-4)i|=|a+(b+2)i|""a^2+(b-4)^2=a^2+(b+2)^2""a^2+b^2-8b+16=a^2+b^2+4b+4""12b=12""b=1"

"a=1-2=-1"

"z=-1+i"